# Enhancing LLM-based Test Generation for Hard-to-Cover Branches via Program Analysis

### v1

@@ IV-C RQ2: Contribution of Each Component in TELPA
To investigate the contribution of each component in TELPA, we compared TELPA with a set of its variants (introduced in Section IV-A3). Table II presents the comparison results in terms of average branch coverage and average line coverage across all projects. As can be seen from Table II, both TELPA and TELPA perform better than their corresponding variants. We further performed a paired Wilcoxon signed-rank test at the significance level of 0.05 to investigate whether TELPA/TELPA significantly outperforms each variant across all projects in terms of branch coverage and line coverage. All p-values are smaller than 0.05, demonstrating the statistically significant contribution of each component in TELPA, regardless of the preceding test generation tools.
@@
In particular, even though we removed one component from TELPA, all these variants still outperform BasicPrompt. This demonstrates the effectiveness of using part of information specific to the challenge posed by hard-to-cover branches, compared to directly integrating as much available information as possible for prompting, which also reflects the negative influence of introducing irrelevant information on LLMs.
@@
We also investigated whether our coverage-based counter-example sampling approach contributes to the overall effectiveness of TELPA by comparing with TELPA that randomly selects counter-examples. Table II shows that TELPA outperforms TELPA regardless of the preceding test generation tools. This result further confirms the importance of using counter-examples even if they are just selected randomly. Moreover, TELPA significantly outperforms TELPA regardless of the preceding test generation tools. For example, TELPA improves TELPA by 9.42% in terms of average branch coverage achieved, while TELPA improves TELPA by 23.07%. The results demonstrate the importance of a diverse set of counter-examples, which can be well achieved by our coverage-based sampling strategy.
IV-D RQ3: TELPA’s Effectiveness with Different Configurations
| Pynguin | CODAMOSA | TELPA | TELPA | TELPA | TELPA | TELPA | |
| Branch | 48.80% | 51.93% | 53.45% | 60.27% | 64.12% | 59.67% | 63.48% |
| Line | 66.02% | 67.83% | 68.38% | 74.16% | 76.36% | 73.34% | 75.51% |
@@ IV-D RQ3: TELPA’s Effectiveness with Different Configurations
We first investigated the effectiveness of TELPA that removes the preceding test generation tool and activates TELPA from scratch. By comparing TELPA with CODAMOSA and Pynguin in Table III, we can see that the TELPA can improve CODAMOSA and Pynguin by 2.92% and 9.53% in terms of average branch coverage and 0.80% and 3.58% in terms of average line coverage. The improvements are not huge, especially compared to the differences between TELPA and CODAMOSA or between TELPA and Pynguin. This is expected, as the time needed for LLMs to generate each test cases is normally much more than that required by SBST techniques. TELPA leverages LLMs for even easy-to-cover branches, preventing it from unleashing its full potential to reach hard-to-cover branches. Instead, CODAMOSA alternately invokes Pynguin and LLMs for test generation, reducing the time needed by LLMs to a large extent, while Pynguin is also efficient as it does not rely on LLMs. Hence, within the same limited testing time, TELPA does not achieve significant coverage improvement over CODAMOSA and Pynguin. This justifies our design choice of adopting preceding test generation tools to balance effectiveness and efficiency.
@@
By examining the data, we found that 603 methods were not covered by CODAMOSA and Pynguin at all within the testing time budget. For these methods, TELPA instead can effectively improve the coverage for 259 of them. On average across these methods, TELPA achieves 21.86% branch coverage and 39.99% line coverage, while TELPA achieves 17.14% branch coverage and 43.80% line coverage. The results further confirm the ability of TELPA in improving the coverage of hard-to-coverage branches.
@@
We then investigated the effectiveness of TELPA and TELPA that employ DeepSeek-Coder-6.7B-instruct [21] instead of Phind-CodeLlama-34B-v2 [9]. From Table III, we found although TELPA/TELPA exhibits small performance reduction compared to TELPA/TELPA, the former still largely outperforms Pynguin/CODAMOSA. For example, TELPA/TELPA improves Pynguin/CODAMOSA by 23.50% and 14.90% in terms of average branch coverage. This demonstrates the generalizability of TELPA under different scales of LLMs to some degree.
IV-E Threats to Validity
@@
The threats mainly lie in parameter settings in TELPA and metrics used. By default, we set the timeframe to two minutes for switching from preceding test generation tools to our LLM-based test generation, the LLM temperature to 1, and the testing time budget to 20 minutes on each software module. Both the timeframe and the testing time budget are longer than the settings in the existing studies [1, 4], demonstrating higher sufficiency of reaching coverage bottlenecks and evaluating test effectiveness. Higher temperature allows TELPA to receive responses from LLMs with higher confidence, which is commonly employed in various tasks [23, 24]. In the future, we will investigate the influence of different settings on the effectiveness of TELPA. Regarding metrics, we have employed the widely-used branch coverage and line coverage, which are also aligned with our goal (i.e., reaching hard-to-cover branches). In the future, we will use more metrics to evaluate generated tests more sufficiently, including naturalness [25] and mutation score [26].
V Discussion
@@
Hallucination. Although TELPA can significantly improve test coverage compared to the state-of-the-art SBST and LLM-based techniques, it is still unable to reach all the hard-to-cover branches within the given testing time budget. Through our analysis, the major reason lies in the hallucination problem, namely LLMs confidently produce incorrect responses. For example, some generated tests by TELPA use non-existent parameters for invoking some methods. Currently, several approaches have been proposed to relieve the hallucination problem in various software engineering tasks, such as fine-tuning LLMs with PPO [27] or DPO [28] algorithms based on high-quality data. In the future, we can incorporate them to further improve the effectiveness of TELPA.
@@
Orthogonality with Prompting Enhancement Methods. TELPA is a novel prompting technique for addressing the challenge of hard-to-cover branches in software testing, which improves the prompt contents specific to this task. Besides improving prompt contents, there are also some other prompting enhancement methods, including various CoT techniques (e.g., SCoT [29] and CCoT [30]) and post-processing techniques (e.g., self-repairing [20]). CoT techniques mainly elicit LLMs to produce intermediate reasoning steps based on the provided prompt for improving LLMs’ effectiveness, while post-processing techniques mainly fix invalid code produced by LLMs based on the given prompt through static analysis or providing error messages to LLMs for self-repairing. They are actually orthogonal to TELPA (improving prompt contents), thus combining them may further improve test coverage.
Generalizability. From our study, the generalizability of TELPA has been demonstrated to a certain extent by evaluating the performance on 27 projects and considering two types of preceding test generation tools (Pynguin and CODAMOSA). Besides, the idea of TELPA is not specific to Python projects (our evaluation benchmark). The involved program analyses in TELPA can be extended to the projects developed in other programming languages, as long as similar data and control flow analyses of this study can be performed on those languages. In fact, TELPA is not restricted to using a certain LLM, and many code generation-related LLMs [31, 32, 33] support program understanding across various programming languages. In the future, we will evaluate TELPA with diverse LLMs on the projects in other programming languages.
VI Related Work
We present traditional and deep-learning-based test generation techniques as related work.

### v2

@@ 4.3. RQ2: Contribution of Each Component in TELPA
To investigate the contribution of each component in TELPA, we compared TELPA with a set of its variants (introduced in Section 4.1.3). Table 3 presents the comparison results in terms of average branch coverage and average line coverage across all the projects. As can be seen from Table 3, both TELPAp and TELPAc perform better than their corresponding variants. We further performed a paired Wilcoxon signed-rank test at the significance level of 0.05 to investigate whether TELPAp/TELPAc significantly outperforms each variant across all projects in terms of branch coverage and line coverage. All p-values are smaller than 0.05, demonstrating the statistically significant contribution of each component in TELPA, regardless of the used preceding test generation tools.
@@
TELPA performs the worst among all variants. This is because the unfiltered invocation sequences can be excessively long and complex, posing significant challenges for LLMs to comprehend the complex sequence, which in turn reduces the effectiveness of test generation. The result demonstrates the contribution of sequence filtering, which could also balance the sequence content and computational burden well. In particular, except TELPA, all other variants still outperform CHATTESTER even when some component is removed from TELPA. This demonstrates the effectiveness of using part of information specific to the challenges posed by hard-to-cover branches, compared to directly integrating unnecessary or insufficient context for prompting, which also reflects the negative influence of introducing irrelevant information on LLMs.
@@
We also investigated whether our coverage-based counter-example sampling approach contributes to the overall effectiveness of TELPA by comparing with TELPA that randomly selects counter-examples. Table 3 shows that TELPA outperforms TELPA regardless of the preceding test generation tools. This result further confirms the importance of using counter-examples even if they are just selected randomly. Moreover, TELPA significantly outperforms TELPA regardless of the preceding test generation tools. For example, TELPAp improves TELPA by 9.01% in terms of average branch coverage achieved, while TELPAc improves TELPA by 11.76%. The results demonstrate the importance of a diverse set of counter-examples, which can be well achieved by our coverage-based sampling strategy. Additionally, TELPA reduced branch coverage by 6.3% on average. The results empirically demonstrate that summarizing the functionality beforehand enhances test generation effectiveness.
4.4. RQ3: TELPA’s Effectiveness with Different Configurations
Cov. | Pynguin | CODAMOSA | TELPA | TELPA | TELPAp | TELPA | TELPAc |
Branch Cov. | 47.71% | 50.87% | 52.75% | 60.34% | 63.98% | 59.79% | 64.06% |
Line Cov. | 64.72% | 67.92% | 68.26% | 73.96% | 75.54% | 74.08% | 74.25% |
@@ 4.4. RQ3: TELPA’s Effectiveness with Different Configurations
TELPA uses existing tools for initial test generation, and utilizes an LLM to generate new tests. In this RQ, we examine how the adoption of existing tools and the choice of the LLM can impact the performance. We first investigated the effectiveness of TELPA that removes the preceding test generation tool and activates TELPA from scratch. By comparing TELPA with CODAMOSA and Pynguin in Table 4, we can see that the TELPA can improve CODAMOSA and Pynguin by 3.70% and 10.56% in terms of average branch coverage and 0.50% and 5.47% in terms of average line coverage. The improvements are not huge, especially compared to the differences between TELPAc and CODAMOSA or between TELPAp and Pynguin. This is expected, as the time needed for LLMs to generate each test cases is normally much more than that required by SBST techniques. TELPA leverages LLMs for even easy-to-cover branches, preventing it from unleashing its full potential to reach hard-to-cover branches. Instead, CODAMOSA alternately invokes Pynguin and LLMs for test generation, reducing the time needed by LLMs to a large extent, while Pynguin is also efficient as it does not rely on LLMs. Hence, within the same limited testing time, TELPA does not achieve significant coverage improvement over CODAMOSA and Pynguin. This justifies our design choice of adopting preceding test generation tools to balance effectiveness and efficiency.
@@
By examining the data, we found that 603 methods were not covered by CODAMOSA and Pynguin at all within the testing time budget. For these methods, TELPA instead can effectively improve the coverage for 259 of them. On average across these methods, TELPAc achieves 21.86% branch coverage and 39.99% line coverage, while TELPAp achieves 17.14% branch coverage and 43.80% line coverage. The results further confirm the ability of TELPA in improving the coverage of hard-to-coverage branches.
@@
We then investigated the effectiveness of TELPA and TELPA that employ DeepSeek-Coder-6.7B-instruct (dee, 2024) instead of Phind-CodeLlama-34B-v2 (phi, 2024). From Table 4, we found although TELPA/TELPA exhibits small performance reduction compared to TELPAp/TELPAc, the former still largely outperforms Pynguin/CODAMOSA. For example, TELPA/TELPA improves Pynguin/CODAMOSA by 26.47% and 17.53% in terms of average branch coverage. To investigate the effectiveness of TELPA when using more advanced LLMs, we evaluated TELPA on GPT-4. We randomly sampled 100 target methods from the benchmark used in previous RQs for evaluation due to the high cost of invoking GPT-4’s APIs. From Table 5, TELPA exhibits similar effectiveness to TELPAp in terms of both average branch coverage and average line coverage on the same 100 sampled methods. This is attributed to the strength of Phind-CodeLlama-34B-v2, a powerful LLM that is specifically trained on code corpus and exhibits close performance to GPT-4 on many code-related tasks (hug, 2024b). The results demonstrate the generalizability of TELPA under different LLMs to some degree.
Cov. | Pynguin | CODAMOSA | TELPA | TELPAp | TELPA |
Branch Cov. | 26.86% | 27.93% | 35.68% | 37.76% | 38.43% |
Line Cov. | 42.13% | 45.75% | 53.66% | 55.97% | 56.08% |
4.5. RQ4: Correctness of Unit Tests Generated by TELPA
Metric | Pynguin | CHATTESTER | TELPAp | CODAMOSA | CHATTESTER | TELPAc |
Execution Pass Rate | 97.44% | 78.57% | 62.48% | 97.66% | 78.82% | 62.07% |
First, we found that all techniques achieve 100% syntax correctness, demonstrating their ability to generate syntactically valid tests without errors. Then, Table 6 compares the execution pass rate of unit tests generated by Pynguin, CODAMOSA, CHATTESTER and TELPA. From the table, both the LLM-based TELPA and CHATTESTER exhibit lower execution pass rates than the search-based Pynguin, indicating the challenge of LLMs in generating semantically valid and executable tests. Note that although CODAMOSA also relies on LLMs, it only leverages LLMs to generate seed tests for Pynguin, and most of the tests are ultimately generated by Pynguin, which results in comparable pass rates between CODAMOSA and Pynguin. TELPA exhibit lower execution pass rates than CHATTESTER. This is because TELPA specifically targets hard-to-cover branches, which often require constructing complex objects and resolving intricate inter-procedural dependencies. These tests are inherently more challenging to execute successfully due to their complexity, involving scenarios prone to runtime errors like incorrect object construction or missing dependencies. In contrast, CHATTESTER generates simpler tests that are less effective at covering complex branches. For example, CHATTESTER’s tests often cover straightforward branches with minimal object construction and fewer inter-procedural dependencies, resulting in higher pass rates but lower overall branch coverage. Specifically, TELPAp and TELPAc cover 21.00% and 21.06% more hard-to-cover branches than CHATTESTER and CHATTESTER, respectively.
We further analyzed the distribution of errors in the generated tests. For TELPA, the tests predominantly fail due to AssertionError (32.0%) and AttributeError (33.5%). AssertionError occurs when the test fails an assertion, which is common in tests targeting complex branches where the expected behavior is harder to predict or verify. AttributeError often happens when the test attempts to access an attribute or method on an object that is not properly constructed or initialized, which is a common issue when dealing with complex object dependencies. Overall, these results confirm that while TELPA generates more complex tests that may have lower execution pass rates, these tests are often essential for achieving higher branch coverage, particularly for hard-to-cover branches involving complex objects and inter-procedural dependencies.
4.6. RQ5: Generalizability of TELPA
Project | EvoSuite | TELPAe |
Chart | 48.64% | 52.45% |
Lang | 85.00% | 88.71% |
Math | 73.57% | 77.51% |
Time | 43.49% | 55.40% |
Average Branch Cov. | 62.68% | 68.52% |
Average Line Cov. | 73.09% | 76.85% |
Branch Type | EvoSuite | TELPAe |
Branches with complex dependencies | 66.53% | 77.42% |
Branches with complex objects | 56.79% | 74.90% |
All hard-to-cover branches (union of the above two types) | 57.55% | 74.59% |
Table 7 presents the comparison results for TELPAe vs. EvoSuite on each project in terms of branch coverage on all branches. The last two rows present the average branch coverage and line coverage across all projects. Table 8 presents the comparison results on hard-to-cover branches. From Table 7 and Table 8, TELPAe achieves consistently higher branch coverage than EvoSuite. Specifically, on average, TELPA improves EvoSuite by 9.32% on all branches and 29.60% on hard-to-cover branches in terms of branch coverage. Note that the improvement on all branches is smaller than the improvement TELPA achieves over Pynguin. One possible explanation is that EvoSuite already incorporates more advanced static analysis techniques than Pynguin, potentially diminishing the additional benefits offered by TELPA. Moreover, other factors, such as differences in programming languages, testing strategies, and even tool implementation details, may also contribute to this discrepancy. Even though, they are still limited in handling intricate inter-procedural dependencies and dynamically computed values, which is confirmed by the larger improvement achieved by TELPA on hard-to-cover branches. The constant pool, for instance, may struggle when required values are not explicitly defined as constants but are derived from complex object construction logic. Similarly, testability transformations excels in cases where flags create coarse fitness landscapes in search-based approaches, making it difficult to find paths leading to high-fitness regions. However, when the program contains complex inter-dependencies or complex object constructions, testability transformation alone may not provide sufficient guidance. The transformation does not inherently handle the semantics of such complex dependencies, which are crucial for generating tests in challenging branches. The results reflects the fact that EvoSuite is already a highly-optimized tool, but TELPA still provides meaningful improvements in terms of code coverage, particularly for those much more challenging branches.
4.7. Threats to Validity
@@
The threats mainly lie in parameter settings in TELPA , metrics and benchmarks used, and the randomness. By default, we set the timeframe to two minutes for switching from preceding test generation tools to our LLM-based test generation, the LLM temperature to 0, and the testing time budget to 20 minutes on each software module. Both the timeframe and the testing time budget are longer than the settings in the existing studies (Lukasczyk and Fraser, 2022; Lemieux et al., 2023), demonstrating higher sufficiency of reaching coverage bottlenecks and evaluating test effectiveness. The results shown in Figure 8 also confirm it. Lower temperature allows TELPA to receive more deterministic responses from LLMs, which is commonly employed in various tasks (Chen et al., 2021; Yan et al., 2025; Tian et al., 2025; Yang et al., 2024a). In the future, we will investigate the influence of different settings on the effectiveness of TELPA. Regarding metrics, we have employed the widely-used branch coverage and line coverage, which are also aligned with our goal (i.e., reaching hard-to-cover branches). We also analyzed the overhead of our static program analysis: on average, the time spent on each module is 1.5 seconds. This overhead is acceptable given its effectiveness. Moreover, this process is performed offline and it could be further accelerated via parallel executions.
Following the existing work (Lemieux et al., 2023), we did not investigate the bug detection effectiveness because test cases must include proper assertions to reveal bugs. However, TELPA focuses on improving branch coverage rather than generating assertions, and various assertion generation approaches (Dinella et al., 2022; Tufano et al., 2022) can be applied to the test cases generated by TELPA. In the future, we will use more metrics to evaluate generated tests more sufficiently.
Regarding benchmarks, one potential threat to the validity of our results is the dynamically typed nature of Python, which presents unique challenges for automated test generation. In dynamically typed languages, type information is not explicitly available at compile time. This makes it difficult for automated test generation tools to infer correct types for object construction and method invocations, potentially leading to ineffective or invalid tests. Furthermore, automated test generation tools for Python, such as Pynguin and CODAMOSA, lack the advanced static analysis techniques used in more mature tools like EvoSuite. This may lead to ineffective or invalid tests and bias the evaluation results. To mitigate this threat, we adapted TELPA to statically typed languages (i.e., Java) and evaluate its effectiveness against EvoSuite on a widely-used Java benchmark Defects4J (Just et al., 2014) in Section 4.6. The results demonstrate that TELPA consistently outperforms EvoSuite on Java projects, which helps mitigating this threat to some extent.
Project | CHATTESTER | TELPAct |
MicroService | 26.47% | 64.71% |
PATool | 89.08% | 90.78% |
DAService | 42.71% | 53.65% |
Average Branch Cov. | 52.75% | 69.71% |
Average Line Cov. | 59.43% | 72.71% |
Furthermore, the projects used for evaluation may appear in the training data of the LLMs used in our study, which can cause the data leakage problem and bias the results (Sallou et al., 2024). Particularly, following the guidelines provided by previous work (Sallou et al., 2024), we also adopted three internal Java projects provided by our industrial partner (Huawei, a global leader company in IT), which can largely reduce the threat from the potential data leakage of open-source projects. The three industrial projects have different functionalities, i.e., a program analysis toolkit, an online micro-service system, and a data analysis framework involving parallel computing and adaptation of design patterns. For ease of presentation, we refer to the three projects as PATool, Microservice and DAService in the following sections. Due to the company policy, we are unable to disclose further details. We constructed TELPAct, which leverages CHATTESTER for preceding test generation and compared it with CHATTESTER. We did not leverage EvoSuite since the three industrial projects use Java 17, but EvoSuite supports up to Java 11.
Table 9 shows the comparison results for TELPAct vs. CHATTESTER on each project in terms of branch coverage. From the table, TELPAct achieves significantly higher branch coverage than CHATTESTER across all three industrial projects. On average, TELPAct improves the branch coverage from 52.75% to 69.71% and the line coverage from 59.43% to 72.71%. The improvements of TELPA on these unseen projects highlight its generalizability and effectiveness in real-world scenarios, independent of potential data leakage from open-source training data.
Regarding randomness, following the previous guideline (Arcuri and Briand, 2011), we repeated all the experiments for 20 times, including experiments on Java and those addressing data leakage threats. With 20 runs, TELPA exhibited a standard deviation of 1.8%, demonstrating consistency. Additionally, we conducted a statistical significance test using the Kruskal–Wallis H test (Kruskal and Wallis, 1952). The result showed no significant differences across runs (p 0.05), indicating stable performance and mitigating this threat to some extent.
5. Case Study
To illustrate the scenarios that TELPA specifically targets and is capable of covering, we present two cases that highlight its ability to address inter-procedural dependencies and complex object construction.
The first case is shown in Figure 10, which demonstrates how TELPA addresses the challenge posed by complex dependencies using branch dependency analysis. The target method format_body processes input content using enabled plugins only if the input matches a specific date pattern validated by is_valid_expression. To cover the branch where plugins are applied, the generated test must provide an input that satisfies is_valid_expression, which requires understanding the regex pattern. Baseline tools, such as Pynguin, generate random inputs (e.g., ”UP/VJeH 5&Db”) that fail the regex check, leaving the target branch uncovered. In contrast, TELPA leverages branch dependency analysis to identify the dependency between format_body and is_valid_expression. By including the source code of is_valid_expression in the prompt, TELPA guides the LLM to infer that valid inputs must adhere to the regex pattern, which requires an underscore in the account part of an email address (e.g., “example_user@gmail.com”). This allows TELPA to generate tests that invoke the plugin branch, while baseline tools fail due to their inability to infer valid inputs. Notably, the test generated by CHATTESTER is omitted here, as it differs from the Pynguin-generated test only in a single random value that still fails the regex check.
The second case is shown in Figure 11, which demonstrates how TELPA addresses the challenge posed by complex object construction using object construction analysis. In this case, the focal method is the schema function, which processes a dataclass (cls) and constructs a schema based on the metadata of its fields. Specifically, it checks if the metadata dictionary of each field contains a key dataclasses_json and uses the associated mm_field to populate the schema. Constructing a valid cls object with the correct metadata structure is critical for covering the branch where metadata.get(mm_field) is not None. Baseline tools struggle with this task: Pynguin and CODAMOSA generate invalid inputs (e.g., integers, booleans) that do not match the expected cls type, while CHATTESTER constructs a valid dataclass but omits the required dataclasses_json metadata. TELPA addresses this challenge by performing object construction analysis to trace the call chain process_schema build_schema schema. It extracts the logic from process_schema, where cls is dynamically instantiated with the required metadata, and incorporates this context into the prompt. As a result, TELPA generates a valid test case with the required metadata structure, ensuring coverage of the target branch.
These cases demonstrate how TELPA leverages program analysis to address challenges proposed by complex dependencies and object constructions. By understanding inter-procedural dependencies and the complex object construction logic, TELPA generates tests that can cover those challenging hard-to-cover branches.
6. Discussion
@@
Hallucination. Although TELPA can significantly improve test coverage compared to the state-of-the-art SBST and LLM-based techniques, it is still unable to reach all the hard-to-cover branches within the given testing time budget. Through our analysis, the major reason lies in the hallucination problem, namely LLMs confidently produce incorrect responses. For example, some generated tests by TELPA use non-existent parameters for invoking some methods. Currently, several approaches have been proposed to relieve the hallucination problem in various software engineering tasks, such as fine-tuning LLMs with PPO (Schulman et al., 2017) or DPO (Rafailov et al., 2024) algorithms based on high-quality data. In the future, we can incorporate them to further improve the effectiveness of TELPA.
@@
Orthogonality with Prompting Enhancement Methods. TELPA is a novel prompting technique for addressing the challenge of hard-to-cover branches in software testing, which improves the prompt contents specific to this task. Besides improving prompt contents, there are also some other prompting enhancement methods, including various CoT techniques (e.g., SCoT (Li et al., 2023) and CCoT (Chia et al., 2023)) and post-processing techniques (e.g., self-repairing (Chen et al., 2024)). CoT techniques mainly elicit LLMs to produce intermediate reasoning steps based on the provided prompt for improving LLMs’ effectiveness, while post-processing techniques mainly fix invalid code produced by LLMs based on the given prompt through static analysis or providing error messages to LLMs for self-repairing. They are actually orthogonal to TELPA (improving prompt contents), thus combining them may further improve test coverage.
Assertion Generation. TELPA focuses on improving branch coverage, a critical precursor to assertion generation. Assertions generated may reveal bugs only if the buggy code is covered by tests. While LLMs are capable of generating complete tests including both test prefixes and assertions (Yuan et al., 2024; Schäfer et al., 2023; Yang et al., 2024b), the quality of these assertions varies. That is, some assertions can effectively determine whether the focal method behaves as expected with specific inputs, while some others are relatively simplistic (e.g., just checking if an object is None). Some works (Dinella et al., 2022; Tufano et al., 2022; Chen et al., 2016) have specifically focused on assertion generation as part of test creation, and these techniques could potentially enhance the quality of LLM-generated assertions. In the future, we plan to refine these assertions by integrating assertion generation techniques and evaluating their effectiveness using metrics like bug detection and mutation scores (Jia and Harman, 2010).
Trade-off in File-Level Analysis. TELPA currently limits its analysis of method invocation sequences to the file hosting the method under test. This design represents a trade-off between comprehensiveness and computational efficiency. The current file typically contains the most relevant usage scenarios for the target method, as it is both the definition location and often the location where it is the most frequently used (Wang et al., 2024b; Yuan et al., 2024). This local context provides sufficient information for generating effective tests in most cases. This approach of focusing on the local context is also widely used in many test generation approaches, such as CHATTESTER (Yuan et al., 2024), ChatUniTest (Chen et al., 2024), and CODAMOSA (Lemieux et al., 2023). While analyzing the entire codebase (e.g., looking at the other classes that might use the target class) could provide more comprehensive results, it would also introduce substantial computational costs. Additionally, cross-file analysis can be particularly challenging in dynamically typed languages like Python, where imports and method invocations are often resolved at runtime. While limiting the analysis to the current file works well in many cases, we acknowledge that it may miss some usage scenarios, particularly when the target method is primarily used in other files. In the future, we can explore project-wide usage patterns to further improve our approach.
7. Related Work